# CReMa: Crisis Response through Computational Identification and Matching of Cross-Lingual Requests and Offers Shared on Social Media

### v1

@@ Abstract
During times of crisis, social media platforms play a vital role in facilitating communication and coordinating resources. Amidst chaos and uncertainty, communities often rely on these platforms to share urgent pleas for help, extend support, and organize relief efforts. However, the sheer volume of conversations during such periods, which can escalate to unprecedented levels, necessitates the automated identification and matching of requests and offers to streamline relief operations. This study addresses the challenge of efficiently identifying and matching assistance requests and offers on social media platforms during emergencies. We propose CReMa (Crisis Response Matcher), a systematic approach that integrates textual, temporal, and spatial features for multi-lingual request-offer matching. By leveraging CrisisTransformers, a set of pre-trained models specific to crises, and a cross-lingual embedding space, our methodology enhances the identification and matching tasks while outperforming strong baselines such as RoBERTa, MPNet, and BERTweet, in classification tasks, and Universal Sentence Encoder, Sentence Transformers in crisis embeddings generation tasks. We introduce a novel multi-lingual dataset that simulates scenarios of help-seeking and offering assistance on social media across the 16 most commonly used languages in Australia. We conduct comprehensive cross-lingual experiments across these 16 languages, also while examining trade-offs between multiple vector search strategies and accuracy. Additionally, we analyze a million-scale geotagged global dataset to comprehend patterns in relation to seeking help and offering assistance on social media. Overall, these contributions advance the field of crisis informatics and provide benchmarks for future research in the area.
@@ I Introduction
    We propose CReMa (Crisis Response Matcher), a systematic approach that incorporates textual, temporal, and spatial features to identify and match requests and offers shared on social media during a crisis.
@@
    We manually curate a new multi-lingual dataset [^1] Dataset will be made available for research purposes upon request. (13k samples) that simulates scenarios of help-seeking and offering assistance on social media.
@@
    We evaluate crisis-specific text classification models, highlighting the superiority of CrisisTransformers over strong baselines like MPNet, RoBERTa, and BERTweet.
  4. 4.
    We show that a cross-lingual embedding space can be effectively used for the extraction of textual features to benefit the matching task. We perform comprehensive cross-lingual experiments across 16 languages.
  5. 5.
    We explore trade-offs between vector search strategies and accuracy, providing insights into the efficiency-accuracy trade-offs in matching tasks.
  6. 6.
    As a case study, we analyze a million-scale geotagged global dataset to explore the distribution of crisis communications concerning requests and offers on social media during the COVID-19 pandemic.
To the best of our knowledge, this study represents the first instances of contributions 2, 4–6 in the crisis informatics domain.
@@ II Related work
For the matching tasks, [3] and [8] rely on a frequency-based approach and pre-trained word2vec, respectively, limiting their scope to the English language. However, we demonstrate later in the paper that embedding approaches such as word2vec, GloVe, fastText, etc., are inadequate for the matching tasks. [8] utilizes POS taggers and NER tools for information extraction. The identification of resources is based on semantic similarity with a predefined list, potentially restricting the methodology to the completeness and relevance of that list and overlooking context-specific terms. Maintaining such lists for multiple languages is impractical. Additionally, their method linearly combines textual similarity and spatial features, resulting in matching of offers even from distant regions. Moreover, temporal aspects are disregarded, leading to a neglect of urgent requests that require immediate attention. In this study, we address the aforementioned issues in the matching tasks by utilizing a cross-lingual embedding space to generate state-of-the-art sentence embeddings while considering both spatial and temporal features. Additionally, unlike prior research, we intentionally increase the difficulty of evaluating matching tasks by employing a manually curated set of hard-coded matched requests and offers.
@@ III-A Regular Expressions
We utilize a set of regular expressions to filter out irrelevant content from the raw data store in order to identify potential requests and offers. We improve upon the initial sets proposed in [3] to develop a more comprehensive set of regular expressions. Initially, a list of regular expressions was manually curated. Subsequently, we matched the texts from datasets provided by [3] and our own help-offer matching datasets (discussed later in Section III-F2) against the regular expressions. We employed a recursive process to add more expressions to incorporate unmatched texts. The final set had 24 regular expressions. Table I presents the top 5 regular expressions utilized for identifying potential requests and offers. These expressions are ranked based on the count of help request and offer texts they successfully match.
@@ III-B1 Fine-tuning
The raw pre-trained models are initially trained for masked language modeling and are designed to be fine-tuned for specific downstream tasks. To fine-tune the pre-trained models for text classification, we added a linear prediction layer to the pooled output. Mean pooling was applied over the token embeddings of an input sequence. Each model was fine-tuned with a maximum of 40 epochs, a batch size of 32, a learning rate of 1e-5, and AdamW as an optimizer with weighted cross-entropy loss. Early stopping was configured with a patience of 8 and a threshold of 0.0001. Fine-tuning was repeated 5 times with different seeds (42, 0, 17, 23, 2024) for each model on 70/10/20 train, validation and test splits. Average performance scores are reported at a 95% confidence interval.
@@ III-C Sentence Encoding
In this study, we utilize the multi-lingual sentence encoder from our earlier work [10] for encoding purposes. The encoder is based on CrisisTransformers’ mono-lingual sentence encoder [9], which underwent further training, as a teacher model in a teacher-student training network on a dataset containing over 128 million multi-lingual parallel sentences, extending its capabilities to a multi-lingual context.
@@ III-F2 Help-Offer Matching Datasets
We used GPT-3.5 Turbo for generating synthetic examples in English to simulate help seeking and offering assistance scenarios for five Australian cities: Sydney, Melbourne, Brisbane, Adelaide, and Perth. The examples were tailored based on probable incidents that have happened in the respective cities. Each example includes the following information: city, help-seeking text language, help-seeking text, help-seeking timestamp, help-seeking location, offer text language, offer text, offer timestamp, and offer location. We manually verified each example for their textual similarity. The initial dataset creation involved the random generation of time and location values. For each example, we considered a maximum 3-day window for timestamps, and a maximum 10-kilometer great-circle distance for locations.
After generating the initial dataset, language translation was performed for each help-seeking text and offer text using the Google Cloud Translate API for 15 commonly used languages in Australia. The following languages were considered: Chinese (Simplified), Chinese (Traditional), Arabic, Vietnamese, Greek, Italian, Hindi, Spanish, Korean, Gujarati, Indonesian, Farsi (Persian), French, German, and Portuguese. This step aimed to simulate multi-lingual social media interactions during a crisis. To diversify the datasets further, we randomly selected request and offer texts in different languages to create additional 10 random datasets (R1–10). Some request and offer examples are provided in Table II. Figure 2 shows language distributions across random datasets 7–10.
Figure 2: Illustrations of language distributions (in %, left: request texts, right: offer texts) across random datasets (R7–10 shown).
@@ IV-B Requests/Offers Matching
Given the abundant resources available for training embedding models in English, we conducted an initial benchmarking of our multi-lingual sentence encoder against several established models including word2vec, GloVe, fastText, InferSent, Universal Sentence Encoder, and Sentence Transformers (SBERT). The matching tasks were done solely based on textual features, i.e., cosine similarity of embeddings. The results, presented in Table VI, show that our CrisisTransformers’ mono-lingual model surpasses all baselines, with the multi-lingual version closely following. Since the multi-lingual model was trained to replicate the embedding space of the mono-lingual model using over 128 million translation pairs, here-on-after, in the study, we employ the multi-lingual model for matching tasks in multi-lingual scenarios, thus establishing a benchmark for future studies.
We compare our approach, referred to as “TTS” (Textual Temporal Spatial), with the baselines “T” and “TS”, where “T” relies on textual features and “TS” relies on linear weighing of textual and spatial features. We experimented with the following proximities wherever applicable: $\delta_{time}=\{30,90,180\}$ (days), $\delta_{location}=\{10,20,30\}$ (kms), and nearest neighbor ($K$) = {25, 50, 100}. Due to the limited size of the synthetic dataset, we conducted an exhaustive search to identify the Top-$n$ nearest offers for each request. This exhaustive search serves as a benchmark for evaluating approximation search strategies. It enables us to investigate the feasibility and accuracy trade-offs of alternative approaches that do not require exploring the entire search space in real-time. Also, the resource classifier is critical for tagging tweets into different categories such as “money’, “volunteer’, etc. This information can be used to limit the search space during matching tasks.
@@
With our approach, the matching tasks exhibited the best performance with $k=100$, $\delta_{time}=30$, and $\delta_{location}=10$, amongst all experiments. Due to space constraints, we present results (in Table VII) from $\delta_{time}=30$ and $\delta_{location}=10$ for all $K$, and make them a reference for discussions. Between the lowest and highest values of $K$, both Top-2 and Top-3 showed higher improvements compared to top-1. This is justified since achieving a perfect match in the Top-1 results is a challenging task, considering the possibility that an offer, perfectly suited for a particular request, might also be an ideal match for other requests due to the random generation of time and location values. We also present results for “TS” in Table VIII; we limit the results due to space constraints to $\delta_{location}=30$, where linear weighing achieved the best results. Except for Gujarati, “TTS” performed better across all languages and random datasets. Next, we discuss the results from our approach in detail.
@@
We observed declining performances as the values of $\delta_{time}$ and $\delta_{location}$ increased. As these values increase, the number of potential offers increase, and finding the one offer hardcoded in the dataset become challenging. When $\delta_{distance}=10$, we observed a decrease in accuracy over time: a 2% decrease on average at $\delta_{time}=90$ and a 4% decrease on average at $\delta_{time}=180$. Similarly, when $\delta_{distance}=20$, we observed a 3% decrease on average at $\delta_{time}=90$ and a 5% decrease on average at $\delta_{time}=180$. $\delta_{distance}=30$ showed similar trend as $\delta_{distance}=20$.
@@
TABLE IX: Statistics of textual similarities where offers corrected matched with “TTS” (Top-3 at K=100).
@@
We examined cases where our approach correctly matched offers in the Top-3 at K=100. For each successful match, we assessed textual similarity. Descriptive statistics of these similarities are presented in Table IX, which includes the mean, quartiles (Q1, Q2, Q3), and maximum values. These statistics provide an overview of similarity scores across various datasets, which assists in understanding the typical range and variability of similarity scores for correctly matched offers using our approach. Results show higher mean similarity for languages such as English (0.43), Italian (0.46), Spanish (0.44), French (0.45), and German (0.45), while the lowest mean accuracies are observed in Hindi (0.33), Farsi (0.33), and Gujarati (0.27). These results are consistent with what we observed regarding matching accuracies across datasets, confirming that textual features play a critical role in the matching tasks as temporal and spatial weights are, in fact, weighing the textual similarity. These results can serve as a reference for future studies to set similarity thresholds across languages for similar matching tasks in crisis informatics.
@@ IV-C Index/search times and accuracy trade-offs
We further examined the trade-offs between index/search times and accuracy of our approach, building upon the findings presented in Table VII, specifically focusing on the English dataset and its Top-3 accuracy. Table X presents a comparison of different search strategies based on their index time, search time, and Top-3 accuracy for various values of $K$. The exhaustive search involves exploring the entire search space; it has the lowest index times (1.42ms on avg.) and highest accuracies (0.85–0.9), resulting in the highest search times (24.52–33.61ms) compared to other methods, which are approximation-based.
@@
TABLE X: Comparison of search methods: index/search times (milliseconds) and Top-3 accuracy trade-offs.

### v2

@@ Abstract
During times of crisis, social media platforms play a crucial role in facilitating communication and coordinating resources. In the midst of chaos and uncertainty, communities often rely on these platforms to share urgent pleas for help, extend support, and organize relief efforts. However, the overwhelming volume of conversations during such periods can escalate to unprecedented levels, necessitating the automated identification and matching of requests and offers to streamline relief operations. Additionally, there is a notable absence of studies conducted in multi-lingual settings, despite the fact that any geographical area can have a diverse linguistic population. Therefore, we propose CReMa (Crisis Response Matcher), a systematic approach that integrates textual, temporal, and spatial features to address the challenges of effectively identifying and matching requests and offers on social media platforms during emergencies. Our approach utilizes a crisis-specific pre-trained model and a multi-lingual embedding space. We emulate human decision-making to compute temporal and spatial features and non-linearly weigh the textual features. The results from our experiments are promising, outperforming strong baselines. Additionally, we introduce a novel multi-lingual dataset simulating help-seeking and offering assistance on social media in 16 languages and conduct comprehensive cross-lingual experiments. Furthermore, we analyze a million-scale geotagged global dataset to understand patterns in seeking help and offering assistance on social media. Overall, these contributions advance the field of crisis informatics and provide benchmarks for future research in the area.
@@ I Introduction
    We propose CReMa (Crisis Response Matcher), a systematic approach that integrates textual, temporal, and spatial features to identify and match requests and offers shared on social media during a crisis.
@@
    We manually curate a new multi-lingual dataset [^1] Dataset will be made available for research purposes upon request. (13k samples) simulating help-seeking and offering assistance scenarios on social media. Utilizing this dataset, we demonstrate that a cross-lingual embedding space can be effectively used to extract textual features for enhancing the matching task.
@@
    As a case study, we analyze a million-scale geotagged global dataset to explore the distribution of crisis communications regarding requests and offers on social media during the COVID-19 pandemic.
@@ II Related work
For the matching tasks, [3] and [8] rely on a frequency-based approach and pre-trained word2vec, respectively, limiting their scope to the English language. However, we demonstrate later in the paper that embedding approaches such as word2vec, GloVe, fastText, etc., are inadequate for the matching tasks. [8] utilizes POS taggers and NER tools for information extraction. The identification of resources is based on semantic similarity with a predefined list, potentially restricting the methodology to the completeness and relevance of that list and overlooking context-specific terms. Maintaining such lists for multiple languages is impractical. Additionally, their method linearly combines textual similarity and spatial features, resulting in the matching of offers even from distant regions. Moreover, temporal aspects are disregarded, leading to a neglect of urgent requests that require immediate attention. In this study, we address the aforementioned issues in the matching tasks by utilizing a cross-lingual embedding space to generate state-of-the-art sentence embeddings while considering both spatial and temporal features. Additionally, unlike prior research, we intentionally increase the difficulty of evaluating matching tasks by employing a manually curated set of hard-coded matched requests and offers.
@@ III-A Regular Expressions
We utilize a set of regular expressions to filter out irrelevant content from the raw data store in order to identify potential requests and offers. We improve upon the initial sets proposed in [3] to develop a more comprehensive set of regular expressions. Initially, a list of regular expressions was manually curated. Subsequently, we matched the texts from datasets provided by [3] and our own help-offer matching datasets (discussed later in Section III-F2) against the regular expressions. We employed a manual recursive process to add more expressions, incorporating unmatched texts until all the texts from both datasets were matched. The final set had 24 regular expressions. Table I presents the top 5 regular expressions utilized for identifying potential requests and offers. These expressions are ranked based on the count of help request and offer texts they successfully match.
@@ III-B1 Fine-tuning
The raw pre-trained models are initially trained for masked language modeling and are designed to be fine-tuned for specific downstream tasks. To fine-tune the pre-trained models for text classification, we added a linear prediction layer to the pooled output. Mean pooling was applied over the token embeddings of an input sequence. Each model was fine-tuned with a maximum of 40 epochs, a batch size of 32, a learning rate of 1e-5, and AdamW as an optimizer with weighted cross-entropy loss. Early stopping was configured with a patience of 8 and a threshold of 0.0001. Fine-tuning was repeated 5 times with different seeds (42, 0, 17, 23, 2024) for each model on 70/10/20 train, validation and test splits. Average performance scores on test data are reported at a 95% confidence interval.
@@ III-C Sentence Encoding
In this study, we utilize the multi-lingual sentence encoder from our earlier work [10] for encoding purposes. The encoder is based on XLM-RoBERTa, which underwent further training as a student in a teacher-student training network with CrisisTransformers’ mono-lingual sentence encoder [9] as the teacher. The network was trained to make XLM-RoBERTa mimic the embedding space of the mono-lingual sentence encoder using a dataset containing over 128 million multi-lingual parallel sentences, including proceedings from the European Parliament, news stories, bilingual dictionaries, translated subtitles, cross-lingual Wiki data, etc., extending the capabilities of the teacher model to a multi-lingual context.
@@ III-F2 Help-Offer Matching Datasets
We used GPT-3.5 Turbo for generating synthetic examples in English to simulate help-seeking and offering assistance scenarios for five Australian cities: Sydney, Melbourne, Brisbane, Adelaide, and Perth. The examples were tailored based on probable incidents that have happened in the respective cities. Each example includes the following information: city, help-seeking text language, help-seeking text, help-seeking timestamp, help-seeking location, offer text language, offer text, offer timestamp, and offer location. We manually verified each example for its textual similarity. The initial dataset creation involved the random generation of time and location values. For each example, we considered a maximum 3-day window for timestamps and a maximum 10-kilometre great-circle distance for locations.
After generating the initial dataset, language translation was performed for each help-seeking text and offer text using the Google Cloud Translate API for 15 commonly used languages in Australia. The following languages were considered: Chinese (Simplified), Chinese (Traditional), Arabic, Vietnamese, Greek, Italian, Hindi, Spanish, Korean, Gujarati, Indonesian, Farsi (Persian), French, German, and Portuguese. This step aimed to simulate multi-lingual social media interactions during a crisis. Altogether, 13,264 samples (16 languages * 829) were curated.
To diversify the datasets further, we created 10 random datasets (R1–10) to introduce variability in the distribution of languages by randomly selecting the language for each help-seeking and offer text.
  • Help-seeking texts: For each sample, a language was randomly selected from the pool of 16 languages. The corresponding help-seeking text in that language was then assigned to the sample.
  • Offer texts: Similarly, a separate random selection process was used to assign a language to the offer text for each sample. This language could be the same or different from the one used for the help-seeking text.
The random selection process ensures that each random dataset (R1–R10) features a unique combination of languages with the same number of samples as 16 homogeneous datasets. This approach results in some datasets having higher concentrations of certain languages, while others might have a more even distribution. This step was necessary to study the presence of some languages in the samples and their effect on overall matching performance. Table II provides some requests and offers examples. Figure 2 shows language distributions across the random datasets.
Figure 2: Illustrations of language distributions (in %, left: request texts, right: offer texts) across random datasets.
@@ IV-B Requests/Offers Matching
Given the abundant resources available for training embedding models in English, we conducted an initial benchmarking of our multi-lingual sentence encoder against several established models, including word2vec, GloVe, fastText, InferSent, Universal Sentence Encoder, and Sentence Transformers (SBERT). The matching tasks were done solely based on textual features, i.e., cosine similarity of embeddings. The results, presented in Table VI, show that our CrisisTransformers’ mono-lingual model surpasses all baselines, with the multi-lingual version closely following. Since the multi-lingual model was trained to replicate the embedding space of the mono-lingual model using over 128 million translation pairs, here-on-after, in the study, we employ the multi-lingual model for matching tasks in multi-lingual scenarios, thus establishing a benchmark for future studies.
We compare our approach, referred to as “TTS” (Textual Temporal Spatial), with the baselines “T” and “TS”, where “T” relies on textual features and “TS” relies on linear weighing of textual and spatial features. We experimented with the following proximities wherever applicable: $\delta_{time}=\{30,90,180\}$ (days), $\delta_{location}=\{10,20,30\}$ (kms), and nearest neighbor ($K$) = {25, 50, 100}. Due to the limited size of the synthetic dataset, we conducted an exhaustive search to identify the Top-$n$ nearest offers for each request. This exhaustive search serves as a benchmark for evaluating approximation search strategies. It enables us to investigate the feasibility and accuracy trade-offs of alternative approaches that do not require exploring the entire search space in real-time. Additionally, the resource classifier can categorize tweets into various classes like “money”, “volunteer”, etc. This classification can help narrow down the search space during matching tasks. However, in this study, we do not use this information to limit the search space due to the small size of the dataset.
@@
TABLE IX: Some examples from the matching tasks (language=$en$, $k$=100, $\delta_{time}=30$ and $\delta_{location}=10$) are provided below. The best-matched offer by the respective method is included.
Request Offer (TTS) Offer (TS) Offer (T)
North Brisbane is in dire need of emergency food and water after the recent disaster. Please help us if you can! To those in need in North Brisbane, I’m prepared to provide emergency food and water. Reach out, and I’ll do my best to assist!
{distance=0.3kms, time=0.06days}
I’m prepared to provide emergency food and water assistance to those in North Brisbane. If you or anyone in the area needs help, please reach out. We’re in this together!
{distance=0.43kms, time=101.9days}
Offering emergency food and water assistance to those affected by the crisis in Melbourne. Your well-being is important. Reach out for support!
{distance=1375.54kms, time=331.8days}
Urgent appeal in Melbourne: Gas leak in the area, and many families are evacuated without warm clothing and blankets. Please help us stay warm! Just got word about the gas leak. I’ve got warm clothing and blankets ready. Let’s support those affected!
{distance=0.55kms, time=0.11days}
Ready to assist with evacuations due to a gas leak in North Melbourne. If you’re in the affected area, reach out for help. Stay safe!
{distance=1.29kms, time=461days}
Ready to assist with evacuations due to a gas leak in North Melbourne. If you’re in the affected area, reach out for help. Stay safe!
{distance=1.29kms, time=461days}
Urgently need a quarantine facility in Adelaide for a family member showing symptoms. Please help! I’m ready to assist Adelaide residents with quarantine facilities. Reach out if you need help!
{distance=0.9kms, time=0.05days}
Adelaide Emergency Services is coordinating quarantine facilities. Contact us for immediate assistance and support.
{distance=0.3kms, time=95.28days}
HealthCare Solutions is offering quarantine facilities in Adelaide. Connect with us for immediate assistance.
{distance=1.5kms, time=176days}
TABLE X: Some mismatched examples with ”TTS” method (language=$en$, $k$=100, $\delta_{time}=30$ and $\delta_{location}=10$).
Request Offer (TTS) True offer
Southeast Brisbane facing a mudslide threat. Need help reinforcing barriers and securing vulnerable areas. Offering my help for cleanup and relief efforts in areas affected by mudslides in Brisbane. Let’s join forces to restore our city. #BrisbaneTogether
{distance=2.67kms, time=2.85days}
I have experience in mudslide prevention. Ready to assist in reinforcing barriers and securing areas at risk!
{distance=2.89kms, time=0.05days}
Emergency evacuation due to a landslide in Adelaide. Need shelter urgently! To those affected by the floods in Adelaide, I’m offering temporary shelter. Stay strong and reach out for assistance!
{distance=2.5kms, time=0.68days}
For those in Adelaide affected by the landslide, I can provide temporary shelter. Stay safe and reach out for support!
{distance=3.35kms, time=0.1days}
In search of a quarantine facility in Melbourne for my family. COVID-19 has hit us hard, and we need a safe space to isolate. Please help! I understand these times have been challenging for many. If anyone in Melbourne needs someone to talk to or needs support, I’m available. #SupportForAll
{not an exact offer}
I’m ready to provide a quarantine space for a family in need. Let’s support each other during these challenging times. Reach out if you require help.
{distance=1.35kms, time=0.06days}
@@
With our approach, the matching tasks exhibited the best performance with $k=100$, $\delta_{time}=30$, and $\delta_{location}=10$, amongst all experiments. Due to space constraints, we present results (in Table VII) from $\delta_{time}=30$ and $\delta_{location}=10$ for all $K$, and make them a reference for discussions. Refer to Table IX for matched examples and Table X for mismatched examples. Between the lowest and highest values of $K$, both Top-2 and Top-3 showed higher improvements compared to Top-1. This is justified since achieving a perfect match in the Top-1 results is a challenging task, considering the possibility that an offer, perfectly suited for a particular request, might also be an ideal match for other requests (refer to first and second examples in Table X) due to the random generation of time and location values. We also present results for “TS” in Table VIII; we limit the results due to space constraints to $\delta_{location}=30$, where linear weighing achieved the best results. Except for Gujarati, “TTS” performed better across all languages and random datasets. Next, we discuss the results of our approach in detail.
@@
We observed declining performances as the values of $\delta_{time}$ and $\delta_{location}$ increased. As these values increase, the number of potential offers increases, and finding the one offer hardcoded in the dataset becomes challenging. When $\delta_{distance}=10$, we observed a decrease in accuracy over time: a 2% decrease on average at $\delta_{time}=90$ and a 4% decrease on average at $\delta_{time}=180$. Similarly, when $\delta_{distance}=20$, we observed a 3% decrease on average at $\delta_{time}=90$ and a 5% decrease on average at $\delta_{time}=180$. $\delta_{distance}=30$ showed similar trend as $\delta_{distance}=20$.
@@
TABLE XI: Statistics of textual similarities where offers corrected matched with “TTS” (Top-3 at K=100).
@@
We examined cases where our approach correctly matched offers in the Top-3 at K=100. For each successful match, we assessed textual similarity. Descriptive statistics of these similarities are presented in Table XI, which includes the mean, quartiles (Q1, Q2, Q3), and maximum values. These statistics provide an overview of similarity scores across various datasets, which assists in understanding the typical range and variability of similarity scores for correctly matched offers using our approach. Results show higher mean similarity for languages such as English (0.43), Italian (0.46), Spanish (0.44), French (0.45), and German (0.45), while the lowest mean accuracies are observed in Hindi (0.33), Farsi (0.33), and Gujarati (0.27). These results are consistent with what we observed regarding matching accuracies across datasets, confirming that textual features play a critical role in the matching tasks as temporal and spatial weights are, in fact, weighing the textual similarity. These results can serve as a reference for future studies to set similarity thresholds across languages for similar matching tasks in crisis informatics.
@@ IV-C Index/search times and accuracy trade-offs
We further examined the trade-offs between index/search times and accuracy of our approach, building upon the findings presented in Table VII, specifically focusing on the English dataset and its Top-3 accuracy. Table XII presents a comparison of different search strategies based on their index time, search time, and Top-3 accuracy for various values of $K$. The exhaustive search involves exploring the entire search space; it has the lowest index times (1.42ms on avg.) and highest accuracies (0.85–0.9), resulting in the highest search times (24.52–33.61ms) compared to other methods, which are approximation-based.
@@
TABLE XII: Comparison of search methods: index/search times (milliseconds) and Top-3 accuracy trade-offs.